# femto-PIXAR: a self-supervised neural network method for reconstructing femtosecond X-ray free electron laser pulses

### v3

@@ V Discussions and Summary
In conclusion, femto-PIXAR is a novel X-FEL diagnostic which we demonstrate by reconstructing the X-ray power profiles from weak X-FEL pulses in a non-standard two-pulse, self-seeding configuration. For this configuration, the classical method of XTCAV analysis, which uses a matching algorithm to subtract similar lasing-off shots for the energy loss calculation, fails to reproduce meaningful correlations with the X-ray gas detector. By contrast, our method is able to distinguish the relative amplitude of each X-ray pulse on a shot-to-shot basis which is necessary for X-ray photon fluctuation spectroscopy experiments and other future experiments at X-FEL facilities. Our approach has various benefits, avoiding the need for energy calibration with an external reference. A neural network based scheme also scales well for the large datasets expected from high-throughput experiments enabled by next generation X-FEL facilities such as the LCLS-II [29]. As opposed to previous deep-learning work, our approach uses self-supervision with a physics-motivated loss function, avoiding the need for high-fidelity simulations. Our method also reconstructs the full 2D phase space, providing interpretability of the network and giving crucial information about when the network might not be applicable for a specific dataset. Future work will consider extensions to more general X-FEL configurations beyond the slotted foil. For example, using diagnostics beyond the XTCAV could enable analysis of configurations with lasing across the full electron bunch [30], e.g. to provide the femtosecond resolution needed for temporal ghost imaging [31]. The analysis chain after reconstructing the lasing-off reference can also make use of machine learning, e.g. incorporating spectral measurements to refine the X-ray pulse profile [32].
VI acknowledgments
Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the US Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. The authors acknowledge support from DESY (Hamburg, Germany). This work has in part been funded by the IVF project InternLabs-0011 (HIR3X). R. Plumley acknowledges support from the US Department of Energy, Office of Science, Basic Energy Sciences, for the Materials Sciences and Engineering Division under contract DE-AC02-76SF00515. J.J.Turner acknowledges support from the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences under Award No. DE-SC0022216.
@@ VII A1: Data preparation
The centered XTCAV images have a shape of 240 x 240 pixel. A median dark image is subtracted and they are normalized to have intensities roughly between 0 and 1 before they are used in the network. Before obtaining the power profiles a median filter is used to identify the signal region and set the other regions to zero. This approach is just used to identify the signal region, no median operation is applied to the image.
VIII A2: Network analysis
@@
We used an U-net architecture as described in [28]. The optimizer is RMSprop of pytorch with a weight decay of 1e-8. Out of some manual tries, we picked a learning rate of 5e-7 and a batch size of 2. We found that we get good reconstructions with this parameter set. A more extensive hyper parameter search might provide slightly better reconstructions or faster conversion of the network. However we found that the data preparation like scaling and centering is way more important than the exact hyper parameters. For the loss we used an MSE loss on the reconstruction, combined with an absolute error on the COM. The COM has a relative low weight of 1e-4 as we found that higher weights significantly worsen the reconstruction.
In all our lasing-on datasets we exclude images with weak XGMD signal ($<5\,\mu$J). With this threshold we discarded about 15% of our data, resulting in about $4\times 10^{5}$ samples with approx. $20$ fs difference, about $1.5\times 10^{6}$ samples with approx. $30$ fs difference and $9\times 10^{4}$ with only one pulse. Weak shots (where essentially no lasing occurs) with $E_{\text{GMD}}<5\,\mu$J are included in the lasing-off dataset to get more training data. This results in about 30K lasing-off samples.
IX A3: Classical analysis
@@
For the classical analysis the signal region is identified as described in Appendix A, and then divided into 120 slices so that all lasing-on and lasing-off data has the same length in slices and lasing-on and lasing-off data can be compared directly. HierarchicalClustering (sklearn.cluster.AgglomerativeClustering) with euclidean distance is used to cluster the current profiles into 500 groups. Normally, a few hundred lasing-offs are used to build a reference set, but to have a fairer comparison all lasing-off images that are used in the training of the network are provided to the algorithm. The COM profile of these groups are averaged and used as reference.
X A3: Implementation of polynomial regression method
@@
For the polynomial regression we used an algorithm described in [33]. Zeng et al. [27] uses a 13 dimensional polynomial, we find that our algorithm does not converge in this case. In Fig. A1 we show one example of the regression method with increasing polynomial degrees. In our benchmark we use the highest degree where the algorithm converges, with a maximum of 13. In Fig. A2 we show resulting regressions for ten different samples. We found that the predictions have a better correlation with the XGMD if we ignore the (obvious wrong) parts of the predicted profile that are negative.
@@
Figure A2: Resulting regression profiles for ten different samples. In case c) and h) the regression predicts additional maxima. In all cases it does not follow the drop completely. Therefore this approach produces a higher spread and an offset compared to the XGMD benchmark.
XI A4: Residual error benchmark
@@
XII A5: Time distribution extracted from reconstructed profiles
@@
Using a double-Gaussian fit over the reconstructed transient X-ray power profiles, we are able to determine the arrival times of the two pulses at the XTCAV target. The time axis in A4 is centered with respect to the COM of the spectrograph in the temporal domain. The reconstructed pulses obtained by the Gaussian fits produce bimodal time distributions that match the X-FEL setups, with average time jitter of 3.3 fs.

### v4

@@ V Discussions and Summary
In conclusion, femto-PIXAR is a novel X-FEL diagnostic method which we demonstrate by reconstructing X-ray power profiles from real X-FEL data. For the configuration used to deliver monochromatic pulse pairs at few fs distance, the classical method of XTCAV analysis, which uses a matching algorithm to subtract similar lasing-off shots for the energy loss calculation, fails to reproduce meaningful correlations with the X-ray gas energy monitor. By contrast, our method is able to distinguish the relative amplitude of each X-ray pulse on a shot-to-shot basis which is necessary for X-ray photon fluctuation spectroscopy experiments [40] and other future experiments at X-FEL facilities that utilize non-standard beam setups. Our approach has various benefits, avoiding the need for energy calibration with an external reference. A neural network based scheme also scales well for the large datasets expected from high-throughput experiments enabled by next generation X-FEL facilities such as the LCLS-II [41]. As opposed to previous deep-learning work, our approach uses self-supervision, avoiding the need for high-fidelity simulations. Our method also reconstructs the full 2D phase space, providing interpretability of the network and giving crucial information about when the network might not be applicable for a specific dataset. Future work will consider extensions to more general X-FEL configurations beyond the slotted foil. For example, using other diagnostics in addition to the XTCAV could enable analysis of configurations with lasing across the full electron bunch [23], e.g. to provide the femtosecond resolution needed for temporal ghost imaging [42] or help overcome hardware limitations that restrict temporal resolution to a few femtoseconds, enabling access to the attosecond regime. The analysis chain after reconstructing the lasing-off reference can also make use of machine learning, e.g. incorporating spectral measurements to refine the X-ray pulse profile [43].
VI Funding
U.S. Department of Energy (DE-AC02-76SF00515., DE-SC0022216); IVF project InternLabs (HIR3X). (0011193).
VII Acknowledgment
Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the US Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. This work has in part been funded by the IVF project InternLabs-0011 (HIR3X). R. Plumley acknowledges support from the US Department of Energy, Office of Science, Basic Energy Sciences, for the Materials Sciences and Engineering Division under contract DE-AC02-76SF00515. J.J.Turner acknowledges support from the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences under Award No. DE-SC0022216. D. Ratner acknowledges support from the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences under Award No. DE-AC02- 76SF00515.
VIII Disclosures
The authors declare no conflicts of interest.
IX Data Availability
The data presented in this study are available on request from the corresponding authors.
X Supplemental Materials
See Supplement 1 for supporting content.
@@ XI A1: Data preparation
The centered XTCAV images have a shape of 240 x 240 pixel. They are centered in both dimensions using the center of mass. A median dark image is subtracted for all the analysis conducted in this paper. Before the data is fed into the network, it is normalized to have intensities roughly between 0 and 1. The reconstruction of the network is re-normalized again to the original range of values to ensure an analogous data handling to the other approaches. Before obtaining the power profiles, a median filter is used to identify the signal region and set the other regions to zero. This approach is just used to identify the signal region, no median operation is applied to the image.
XII A2: Obtaining the X-ray power profile from the XTCAV spectrograph
The time-dependent power profile $P(t)$ of the X-ray pulse can be derived by the following expressions from FEL theory [30, 31].
$P(t)=\frac{I(t)}{e}(\langle E_{\text{off}}(t)\rangle-\langle E_{\text{on}}(t)\rangle),$ (S1)
and
$P(t)\propto I(t)^{2/3}(\sigma^{2}_{\text{on}}(t)-\sigma^{2}_{\text{off}}(t)),$ (S2)
where $I(t)$ is the beam current, $e^{-}$ the electron charge, $\langle E_{\text{off}}(t)\rangle$ and $\sigma_{\text{off}}(t)$ are the time-dependent $e^{-}$ energy expectation value and variance for the case when the energy distribution of the electrons in the electron bunch was not influenced by lasing, and $\langle E_{\text{on}}(t)\rangle$ and $\sigma_{\text{on}}(t)$ for the case when the electron distribution was shaped by the lasing process. (S2) can only give proportionality and requires power normalization with the GEM. As this is the value we want to benchmark with, we decided to use equation (S1) in order to obtain the power profiles.
XIII A3: Network analysis
@@
We used an U-net architecture as described in [35]. The optimizer is RMSprop of pytorch with a weight decay of 1e-8. Based on a few manual trials, we chose a learning rate of $1\times 10^{-4}$ and a batch size of 32. A more extensive hyper-parameter search might yield slightly better reconstructions or faster convergence.
For the loss function $L$, we used a weighted sum comprising the mean-squared error between the reconstruction image $r$ and the input image $i$, both containing $240^{2}$ pixels, combined with the mean absolute error of the centers of mass (COM), denoted by $com_{i}$ and $com_{r}$:
$L=\frac{1}{240^{2}}\sum_{j}(i_{j}-r_{j})^{2}+w_{COM}\,\langle|com_{i}-com_{r}|\rangle$ (S3)
The weight of the COM term, $w_{COM}$, was set to the relatively low value of $1\times 10^{-4}$, as higher weights significantly degraded the reconstruction.
X-FEL shots where essentially no lasing occurs with $E_{\text{GEM}}<5\,\mu$J are included in the lasing-off dataset to get more training data. This results in about 30k lasing-off samples that are used for training, with an additional 1.5k frames being used as validation dataset to observe for overfitting. Additionally, 460 lasing-off samples were held out as a test set; they were not exposed to the network during training and were used only for the residual-error benchmark. In total, our lasing-on dataset consists of about $4\times 10^{5}$ samples with approximately 202020 fs difference, about $1.5\times 10^{6}$ samples with approximately 303030 fs difference and $9\times 10^{4}$ with only one pulse. Training was stopped when the validation loss failed to decrease by at least $7.5\times 10^{-7}$ within a patience window of 25 steps. On an RTX-2080 S GPU this required roughly 3 hours without any specific optimization. We do not expect a network trained on one dataset to generalize well enough to handle a dataset originating from a completely new FEL setup. However, training is sufficiently fast that retraining a network for a new setup should be straightforward.
XIV A4: Masking procedure
During training, masks are placed at random positions to teach the network to reconstruct missing regions at various positions. Because we wish to mask only regions where the electron-bunch signal is present, we restrict the masking area (to pixel 60–180, see Fig. A1). We further constrain the random masks so that the minimum distance between them is 0 (i.e. they do not overlap) and the maximum distance is 40 pixels. The mask size is fixed at 16 pixels; however, the exact size is not critical for the reconstruction. A discussion of the influence of mask size on the predicted pulse profiles and pulse separation follows below.
During inference (i.e. when the trained network is used for analysis), the masks must be positioned so that they cover the lasing regions. We let the network determine these positions directly. To accelerate the mask search, we do not search the entire 240-pixel frame but instead define a range where the lasing feature is expected. We explicitly exclude the borders of the electron bunch, because shot-to-shot variations are greatest there and cannot be reliably inferred from the unmasked part. This masking area extends from pixel 100 to pixel 161 (see Fig. A1). We then place one mask at a dummy position outside the expected lasing region (pixel 62). This was a position that was also used during network training. This step is necessary because the network was always trained with two masks; as expected, artifacts appear in the reconstruction if only a single mask is used.
Figure A1: The red lines mark the borders of the possible masking region during inference, and the black line marks the left border of the dummy mask, which is intentionally placed outside the lasing region during the search for the correct mask positions.
We then construct a set of images, each containing the dummy mask plus a candidate position for the second mask, with every possible candidate position included in the set. The NN then processes the entire set. We identify the input mask position that produces the largest energy difference (mask A) and the position with the second-largest difference that is at least one mask width away (mask B), ensuring the masks do not overlap. For further evaluation, the mask positions are ordered so that mask1 is the leftmost of mask A and mask B. These two masks are then fed into the network to obtain the final reconstruction.
XIV.1 A5: Influence of the mask size on GEM correlation and time distribution
We chose a mask size of 16 pixels based on the extent required to cover the entire lasing feature of a strong photon pulse. To assess the influence of mask size, we trained and evaluated four additional networks identical in all respects except for the mask size, which was set to 8, 12, 20, and 24 pixels, respectively. The results on the 20 fs separation dataset are shown in Fig. A2 and Fig. A3.
Fig. A2 shows that with a mask size of 8 pixels, the predicted photon pulse energies are lower because not all signal is captured. For a mask size of 24 pixels, the distribution is noticeably broader due to the inclusion of a wider noise background around the lasing region. However, near 16 pixels the exact mask size appears quite robust and does not substantially affect the correlation.
Similarly, the predicted difference in arrival time is not strongly influenced by mask size. Only for the largest mask size (24 pixels) is a significantly higher arrival-time difference predicted.
A larger mask size increases the risk that inaccuracies in the lasing-off electron-bunch energy distribution outweighs the actual lasing signal, leading to an inaccurate pulse prediction.
In summary, it is very straight-forward to find a matching mask size by inspection of the strongest pulses and does not require precise tuning for accurate reconstructions. We found that our networks results are robust against different mask sizes as long as they are in a reasonable range.
Figure A2: Plots of GEM correlations for different mask sizes. The correlation is stable around a mask size of 16 pixels. Each 2D histogram is normalized independently so that its maximum value is 1.
Figure A3: Distribution of the difference in pulse arrival times for both two-pulse setups and various mask sizes.
XV A6: Classical analysis
@@
For the classical analysis the signal region is identified as described in the chapter data preparation, and then divided into 120 slices so that all lasing-on and lasing-off data has the same length in slices and can be compared directly. This method of finding the signal region and slicing also solves the challenge of aligning lasing on and lasing off horizontally (in the time domain), something that is necessary due to timing jitter. HierarchicalClustering with euclidean distance is used to cluster the projection on the y-axis (the time axis), into 500 groups. The projection on the y-axis corresponds to the electronic current if the signal is normalized with the total charge in the electron bunch, measured by an independent detector at the beam dump. Normally, a few hundred lasing-offs are used to build a reference set, but to have a fairer comparison all lasing-off images that are used in the training of the network are provided to the algorithm. The COM profiles of each of these groups are averaged and used as reference.
In order to find the best matching reference group, the $y$-projections of a respective shot is compared to each of the $y$-projections of the lasing-off clusters. This is done with the Pearson product-moment correlation coefficients. The non-diagonal elements of this matrix form an array where each element is the correlation coefficient between the lasing on current profile and each lasing off current profiles. This gives a measure how strongly the fixed lasing on data point and each lasing off reference current profile are related. The cluster with the strongest correlation is taken as the lasing off reference for this specific data point.
XVI A7: Implementation of polynomial regression method
@@
For the polynomial regression we used an algorithm described in [44]. Zeng et al. [34] uses a 13 dimensional polynomial, we find that our algorithm does not converge in this case. In Fig. A4 we show one example of the regression method with increasing polynomial degrees. In our benchmark we use the highest degree where the algorithm converges, with a maximum of 13. In Fig. A5 we show resulting regressions for ten different samples. We found that the predictions have a better correlation with the GEM if we ignore the (obviously wrong) parts of the predicted profile that are negative. Still, this method struggles to predict the detailed structure within the masked region, greatly overestimating the influence of the lasing process.
@@
Figure A5: Resulting regression profiles for ten different samples. In case c) and h) the regression predicts additional maxima. In all cases it does not follow the drop completely. Therefore this approach produces a higher spread and an offset compared to the GEM benchmark.
XVII A8: Residual error benchmark
An additional method to benchmark our method is to evaluate the generated power profiles for data where no lasing process took place, as we have a ’ground truth’, meaning the actual value available. Ideally, for this setup, the difference between the XTCAV data and the reconstruction would be zero, resulting in a mean squared error of zero, and also the integral of the power profiles would be zero. For this benchmark we used 460 lasing-off samples that were not part of the network training (not in the training or validation set) and used each of them 10 times with randomly set masks.
Fig. A6 shows the histogram of relative squared errors for the lasing off reconstructions, evaluated within the masked regions. To demonstrate that these values indeed indicate strong similarity, we also show the histogram of the relative squared difference between the lasing off shots and their respective subsequent shots - where the difference is often not clearly visible by the eye. We show that our approach clearly succeeds in this benchmark. The relative squared error is calculated using the following formula:
$\text{relative squared error}=\frac{\sum_{x,y}(r_{x,y}-i_{x,y})^{2}}{\sum_{x,y}i},$ (S4)
where $r$ and $i$ are once again the reconstructed and input images.
Figure A6: The left image shows an example of the mean error between a lasing-off shot and its reconstruction using the femto-Pixar approach. The middle image shows, for comparison, the difference between a lasing off shot and its subsequent shot. The histogram displays the mean squared error for all data points from the two approaches.
@@ XVII A8: Residual error benchmark
XVIII A9: Time distribution extracted from reconstructed profiles
@@
[…]